OpenStudy (josee):

When f is defined by f(x) = sqrt(x) , find a so f'(a) is two times the value of f'(4).

OpenStudy (sriram):

f'x=1/2sqrtx 1/2sqrta=2*1/2sqrt4 hence a=a=1

OpenStudy (anonymous):